# Optimum Spectrum Extension for PAPR Reduction of DFT-s-OFDM

### v1

@@ I Introduction
Moreover, a common misconception is that larger SE always reduces the PAPR, with its size constrained only by bandwidth-efficiency concerns. In contrast, this paper draws on new insights into FDSS-SE design, showing that for a given FDSS window there exists both an optimal SE size for PAPR minimization – beyond which PAPR performance degrades, and an optimal SE size for rate maximization – below which rate decreases. Hence, if in 6G the FDSS windows remain proprietary and UE-specific as in 5G, i.e., implemented by manufacturers and not optimizable by the network, then enforcing a single SE size for all UEs may be suboptimal and even detrimental for some of them.
@@
Next, we examine the spectral efficiency of FDSS-SE. While conventional wisdom would suggest that SE always reduces rate, we derive a generic capacity expression for DFT-s-OFDM with FDSS-SE and the combining receiver of [10]. Our generic FDSS-SE approach simplifies the prior formulation of this receiver, and the capacity result shows that there is another SE size for maximizing the rate, dependent of the FDSS window and independent of the bandwidth, but, unlike the PAPR-optimal SE size, dependent also of the SNR. The PAPR-optimal and rate-optimal SE sizes differ, but a balanced choice can improve both metrics simultaneously.
@@ II-A DFT-s-OFDM with FDSS and Spectrum Extension
A DFT-s-OFDM symbol with cyclic prefix (CP) is defined for samples $N_{\rm cp}\leq n\leq N_{\rm fft}-1$ as
@@
II-B Pulse Time-Multiplexing Intepretation
@@
$p_{0}[n]=\frac{e^{{\rm j}\frac{\pi}{N_{\rm fft}}n(N_{\rm sc}-1)}}{\sqrt{N_{\rm data}}}\frac{\sin\left(\pi\frac{N_{\rm sc}}{N_{\rm fft}}n\right)}{\sin\left(\frac{1}{N_{\rm fft}}n\right)}.$ (8)
@@ II-C FDSS Windows
With $\beta=0$, we recover the classical Hann window up to normalization, while with $\beta=1$ corresponds to no FDSS shaping. This window is derived based on 3-tap filters commonly-used by industry in 3GPP, see Appendix -A. In this convenient form, the maximum power ripple of this FDSS window is tunable and directly given by its shaping parameter as $\approx 20\log_{10}\beta$. Moreover, the FDSS windows from 3-tap filters are not exactly symmetric and we applied an half index shift to correct this, which has a noticeable effect only for very narrow bands. Specifically, the specific 3-tap filter used in [22, 23, 24] are equivalently obtained as $W_{\beta}[k-1/2]$ using (9) with $\beta=-11$ and 14-14 dB, respectively.
@@ III PAPR: Definition and Relevance
${\rm PAPR}=\frac{\max_{0\leq n<N_{\rm fft}}|s[n]|^{2}}{\mathbb{E}\left[|s[n]|^{2}\right]}.$ (11)
where $\mathbb{E}\left[\cdot\right]$ is the statistical expectation. In numerical evaluations, we will use $N_{\rm fft}=2048$. PAPR of the digital OFDM signal provides an accurate PAPR estimation of the corresponding transmitted analog OFDM signal when oversampling is at least $N_{\rm fft}/N_{\rm data}=4$ [25].
@@ III-A Statistical Mean Power versus Instantaneous Mean Power
The PAPR in (11) is computed using the statistical mean power (i.e., the expectation in the denominator) as in many academic works, e.g., [14, 25, 4]. Alternatively, several other works define the PAPR using the instantaneous arithmetic mean power over a considered signal lengh $N$, i.e. ${\rm PAPR}^{\prime}=\frac{{\max_{0\leq n<N}}|s[n]|^{2}}{\frac{1}{N}\sum_{n=0}^{N-1}{|s[n]|^{2}}}$, and for this, often the signal length is one OFDM symbol as, e.g., in [11].
@@
Fig. 4 illustrates these considerations with $N_{\rm sc}=24$ and deformed Hann window, where the $10^{-3}$ complementary cumulative distribution function (CCDF) level of the PAPR is plotted along increasing FDSS shaping attenuation. The PAPR is computed for both a single OFDM symbol and a 14 CP-OFDM symbol signal [^2] The maximum in the numerator of (11) is applied on the corresponding signal, i.e. for 14 OFDM symbols the peak is taken over a larger sample set, and so its PAPR is larger than with one OFDM symbol.. In the absence of FDSS (0 dB), both PAPR definitions yield the same value for $\pi/2$-BPSK and QPSK. However, as FDSS shaping increases, the two PAPR definitions diverge. This difference is small for $\pi/2$-BPSK, and for QPSK when signals comprise 14 CP-OFDM symbols. Similarly, for a single OFDM symbol with a larger number of subcarriers, the PAPR based on instantaneous mean power would approach that based on statistical mean power. However, as shown, for the case of QPSK with FDSS, the PAPR computed using the instantaneous mean power is significantly higher for single narrowband OFDM symbol than other cases, and thus fails to consistently reflect performance trends.
@@ III-B Linear Fitting of Power De-rating by PAPR and CM
The PAPR measures the envelope fluctuation of a signal. Peaks in signal envelope either cause distortion by entering the nonlinear region of the power amplifier; or alternatively require power de-rating to avoid such distortion which directly reduces the link SNR. 3GPP standards specify a maximum power reduction (MPR) level that limits the allowed de-rating (also referred to as output back-off or offset) to ensure conformance with RF requirements, including those for signal distortion and adjacent channel interference. In practical systems, the required power de-rating depends on the bandwidth allocation size and its position within the channel bandwidth, typically being higher for either large allocations or small allocations located near the channel edge. Thus, in the corresponding 3GPP specification [26], edge allocations have larger MPR values compared to inner allocations, and larger allocated bandwidth carry additional MPR margins. In general, higher modulation orders lead to higher PAPR and thus allow greater MPR values.
@@
The cubic metric (CM) is an alternative to PAPR that has been considered superior in 3GPP for quantifying MPR. The (raw) cubic metric is defined as ${\rm CM}\,{\rm[dB]}=20\log_{10}{\rm CM}$ with ${\rm CM}={\rm rms}\left(\left|\frac{s[n]}{{\rm rms}(s[n])}\right|^{3}\right)$ where ${\rm rms}\left(s[n]\right)=\sqrt{\sum_{n=0}^{N-1}|s[n]|^{2}}$. The metric was introduced during late 3G UMTS standardization, based on nonlinear amplifier modeling, with the argument that it more accurately predicts power de-rating than PAPR [27]. From its inception, CM has been paired with linear regression coefficients to match MPR measurements. In fact, the 3G UMTS specification [28] includes multiple shifted, scaled, and quantized versions of CM depending on the system configuration. For 4G LTE, the CM fitting was updated [29] based on OFDM signal measurements, and different regression coefficients were recommended for two different bandwidth configurations. Since then, 5G NR has reused these definitions from [27], originally derived for 4G – and in some cases from 3G work, see e.g., [30]. However, applying the early-4G linear fit often leads to negative power de-rating estimates when evaluating modern low-PAPR waveforms, such as DFT-s-OFDM with $\pi/2$-BPSK and FDSS as well as other schemes see e.g. [31], which raises questions about the practical meaning of these CM-based MPR estimate.
@@
It is worth emphasizing that CM may provide much better power de-rating estimation than PAPR, primarily because it is always used with fitted regression coefficients from measurements, whereas PAPR is typically evaluated in its raw form, without linear fitting. However, nothing prevents applying a linear regression to PAPR as well, to estimate more realistic de-rating values. In fact, in the original 3GPP contribution introducing CM [27], a linear fitting of PAPR was also considered: ${\rm MPR}\approx 0.855({\rm PAPR}-3)$. Although CM provided a slightly better overall fit, the relationship between power de-rating and PAPR still followed a linear correlation reasonably well. More recent evaluations, such as [32], confirms this.
In any case, neither CM nor PAPR can fully predict realistic power de-rating values. For instance, in [24], both the CM and PAPR of QPSK with FDSS appear worse than with clipping, while the measured power de-rating is actually better. Most importantly, the impact of bandwidth allocation size and position relative to the system bandwidth is not captured by either CM or PAPR.
@@ IV-A PAPR Bounds
Moreover, for regular QAM, note that the maximum symbol power $A_{\mathcal{C}}^{2}$ occurs only for a subset of constellation points at the four edges , forming a QPSK sub-constellation with larger amplitude. We can therefore expect that symbol combination of the this sub-QPSK constellation will yield some of the highest signal peaks, leading to:
${\rm PAPR}^{\rm U}(\mathcal{C}_{\rm QAM},N_{\rm e},L)\approx A_{\mathcal{C}_{\rm QAM}}^{2}\times{\rm PAPR}^{\rm U}(\mathcal{C}_{\rm QPSK},N_{\rm e},L)$ (16)
@@ IV-B Best Shifting Value
In Fig. 6, we show the deviation of the $10^{-3}$-CCDF PAPR as a function of the SE ratio $N_{\rm e}/N_{\rm sc}$ for QPSK with $N_{\rm sc}=96$ and Kaiser FDSS window with $\kappa=2$, which corresponds to a maximum power attenuation of $7.15$ dB. The PAPR deviation between i) $L$ from (18) corresponding to the symmetric extension in [10], and ii) $L$ from (19), is computed with reference to iii) $L=0$, corresponding to the single-side extension in [14]. The figure also show a prediction of this PAPR deviation using the bound (13) with the corresponding $L$ values. As seen, for QPSK the symmetric extension from [10] is always worse than the single-side extension from [14]. Similar results can be obtained for higher-order QAMs. The maximum PAPR gain over [10] by using a different $L$ is only a fraction of a dB, here up to 0.5 dB, and this gain decreases as the SE size increases.
@@ IV-C Optimum SE Size $N_{\rm e}^{({\rm min\,PAPR})}$ for Minimum PAPR
IV-C1 Existence of an Optimal SE
It is well known in the literature that SE can further reduce the PAPR on top of FDSS. However, it less recognized that there exists an optimum SE size that minimizes the PAPR, and that a very large SE may even increase it.
To intuitively understand this, consider the pulse-shaping filters in (6), derived from the band-limited FDSS window $W[k]$ in (7). Their amplitude essentially follows a sinc shape, with more or less attenuated side lobes. The kernel filter shape in (7) is independent of $N_{\rm e}$, except for a scaling factor $\frac{1}{N_{\rm sc}-N_{\rm e}}$, which causes the peak energy of each pulse to increase as $N_{\rm e}$ increases. In parallel, since there are $(N_{\rm sc}-N_{\rm e})$ pulses per OFDM symbol, the pulses are spaced $\frac{N_{\rm fft}}{N_{\rm sc}-N_{\rm e}}$ samples apart in (6). This spacing increases with $N_{\rm e}$, which in turn changes the degree of pulse overlap due the sinusoidal shape of the pulses. Both effects combined, depending of $N_{\rm e}$, their overall combining may decrease or increase as a function of $N_{\rm e}$, leading to a non-monotonic PAPR variation.
@@ IV-C1 Existence of an Optimal SE
In Fig. 8, we plot several examples of PAPR versus the SE ratio $N_{\rm e}/N_{\rm sc}$. Each plot shows the maximum PAPR from $10^{6}$ Monte Carlo points, the $10^{-1}$, $10^{-2}$, and $10^{-3}$ CCDF, and the bounds from (13) and (15). All subplots is with QPSK, except Fig. 8(d) with $\pi/2$-BPSK and Fig. 8(e) with 16QAM. All have $N_{\rm sc}=96$ except Fig. 8(f) with $N_{\rm sc}=48$. The first row, i.e. Figs. 8(a)–(c), uses different FDSS windows, while the second row, i.e. Figs. 8(d)–(f), varies the modulation or bandwidth, all with a Kaiser FDSS window ($\kappa=2$). Shift values match single-side extension ($L=0$) in the literature, except Fig. 8(d) where $L$ is chosen according to (18).
@@
Fig. 8 clearly shows that the PAPR does not decrease monotonically with $N_{\rm e}$. In all cases, the PAPR decreases up to an optimum SE size $N_{\rm e}^{({\rm min\,PAPR})}$, beyond which it increases again. Selecting $N_{\rm e}>N_{\rm e}^{({\rm min\,PAPR})}$ is detrimental not only for spectral efficiency but also for PAPR reduction. We have systematically verified the existence of this global optimum across many scenarios. Remark that even without FDSS [Fig. 8(a)], SE can provide some PAPR reduction and so $N_{\rm e}^{({\rm min\,PAPR})}>0$.
@@
From Figs. 8(a)–(c), one can observe that $N_{\rm e}^{({\rm min\,PAPR})}$ changes with the FDSS window, while the PAPR curves are almost identical between QPSK [Fig. 8(b)] and 16-QAM [Fig. 8(e)]. Similarly, the PAPR behavior is consistent for QPSK with different bandwidths [Figs. 8(b) and 8(f)]. In [15], we showed by tedious simulations that $N_{\rm e}^{({\rm min\,PAPR})}$ as a percentage of $N_{\rm sc}$, remains roughly constant for different $N_{\rm sc}$, given different examples of FDSS windows. In the following, we show that PAPR bounds can provide a good characterization of this optimum without exhaustive simulations.
IV-C2 Characterizing $N_{\rm e}^{({\rm min\,PAPR})}$

### v2

@@ I Introduction
Moreover, while it is well established that SE can enable further PAPR reduction on top of FDSS, it is less recognized that an excessively large SE may instead increase it. The prevailing view being that SE size should be limited, but solely by bandwidth-efficiency considerations. Conversely, if ignoring the PAPR effect, SE is commonly regarded as being inherently detrimental to spectral efficiency. In contrast, this paper draws on new insights into FDSS-SE design, showing that for a given FDSS window there exists both an optimal SE size for PAPR minimization – beyond which PAPR performance degrades, and an optimal SE size for rate maximization – below which rate decreases. Hence, if in 6G the FDSS windows remain proprietary and UE-specific as in 5G, i.e., implemented by manufacturers and not optimizable by the network, then enforcing a single SE size for all UEs may be suboptimal and even detrimental for some of them.
@@
Next, we examine the spectral efficiency of FDSS-SE. While conventional wisdom would suggest that SE always reduces rate, we derive a generic capacity expression for DFT-s-OFDM with FDSS-SE and the combining receiver of [31]. Our generic FDSS-SE approach simplifies the prior formulation of this receiver, and the capacity result shows that there is another SE size for maximizing the rate, dependent of the FDSS window and nearly invariant of the bandwidth, but, unlike the PAPR-optimal SE size, dependent also of the SNR.
For both metrics, PAPR and rate, analysis and results are supported by a semi-analytical approach based on numerical optimization of theoretical bounds. The PAPR-optimal and rate-optimal SE sizes are in general different, providing a bounded range of relevant SE sizes within which a balanced choice can improve both metrics simultaneously.
@@ II-A DFT-s-OFDM with FDSS and Spectrum Extension
Figure 1: Block diagram of DFT-s-OFDM with FDSS-SE.
The DFT-s-OFDM transmitter with FDSS-SE is schematized in Fig. 1 A DFT-s-OFDM symbol with cyclic prefix (CP) is defined for samples $N_{\rm cp}\leq n\leq N_{\rm fft}-1$ as
@@
II-B Pulse Time-Multiplexing Interpretation
@@
$p_{0}[n]=\frac{e^{{\rm j}\frac{\pi}{N_{\rm fft}}n(N_{\rm sc}-1)}}{\sqrt{N_{\rm data}}}\frac{\sin\left(\pi\frac{N_{\rm sc}}{N_{\rm fft}}n\right)}{\sin\left(\pi\frac{1}{N_{\rm fft}}n\right)},$ (8)
@@ II-C FDSS Windows
With $\beta=0$, we recover the classical Hann window up to normalization, while with $\beta=1$ it corresponds to no FDSS shaping. This window is derived based on 3-tap filters commonly-used by industry in 3GPP, see Appendix -A. In this convenient form, the maximum power ripple of this FDSS window is tunable and directly given by its shaping parameter as $\approx 20\log_{10}\beta$. Moreover, the FDSS windows from 3-tap filters are not exactly symmetric and we applied an half index shift to correct this, which has a noticeable effect only for very narrow bands. Specifically, the specific 3-tap filter used in [7, 10, 8] are equivalently obtained as $W_{\beta}[k-1/2]$ using (9) with $\beta=-11$ and 14-14 dB, respectively.
@@ III PAPR: Definition and Relevance
${\rm PAPR}=\frac{\max_{0\leq n<N_{\rm fft}}|s[n]|^{2}}{\mathbb{E}\left[|s[n]|^{2}\right]}$ (11)
where $\mathbb{E}\left[\cdot\right]$ is the statistical expectation. Otherwise stated, $N_{\rm fft}=2048$ in numerical evalutations. PAPR of the digital OFDM signal provides an accurate PAPR estimation of the corresponding transmitted analog OFDM signal when oversampling is at least $N_{\rm fft}/N_{\rm data}=4$ [22].
@@ III-A Statistical Mean Power versus Instantaneous Mean Power
The PAPR in (11) is computed using the statistical mean power (i.e., the expectation in the denominator) as in many academic works, e.g., [35, 22, 26]. Alternatively, several other works define the PAPR using the instantaneous arithmetic mean power over a considered signal lengh $N$, i.e. ${\rm PAPR}^{\prime}=\frac{{\max_{0\leq n<N}}|s[n]|^{2}}{\frac{1}{N}\sum_{n=0}^{N-1}{|s[n]|^{2}}}$, and for this, often the signal length is one OFDM symbol as, e.g., in [20, 29, 23, 28, 21, 19].
@@
Fig. 5 illustrates these considerations with $N_{\rm sc}=24$ and deformed Hann window, where the $10^{-3}$ complementary cumulative distribution function (CCDF) level of the PAPR is plotted along increasing FDSS shaping attenuation. The PAPR is computed for both a single OFDM symbol (1OS) and a 14 CP-OFDM symbol signal [^2] The maximum in the numerator of (11) is applied on the corresponding signal, i.e. for 14 OFDM symbols the peak is taken over a larger sample set, and so its PAPR is larger than with one OFDM symbol. (14OS). In the absence of FDSS (0 dB), both PAPR definitions yield the same value for $\pi/2$-BPSK and QPSK. However, as FDSS shaping increases, the two PAPR definitions diverge. This difference is small for $\pi/2$-BPSK, and for QPSK when signals comprise 14 CP-OFDM symbols. Similarly, for a single OFDM symbol with a larger number of subcarriers, the PAPR based on instantaneous mean power would approach that based on statistical mean power. However, as shown, for the case of QPSK with FDSS, the PAPR computed using the instantaneous mean power is significantly higher for single narrowband OFDM symbol than other cases, and thus fails to consistently reflect performance trends.
@@ III-B Linear Fitting of Power De-rating by PAPR and CM
The PAPR quantifies the envelope fluctuation of a signal. Large envelope peaks either drive the power amplifier into its nonlinear region or require transmit power back-off to avoid distortion, which directly reduces the link SNR. In 3GPP systems, this back-off is constrained by a maximum power reduction (MPR), defined to ensure compliance with RF requirements such as distortion and adjacent-channel interference. In practice, the required MPR depends on the allocated bandwidth size and its position within the channel band, with larger allocations or edge allocations typically requiring larger MPR margins. Higher-order modulations generally exhibit higher PAPR and therefore incur larger MPR.
@@
The cubic metric (CM) has been introduced during late 3G UMTS standardization as an alternative to PAPR for estimating power de-rating, motivated by nonlinear amplifier modeling [13]. From its inception, CM has always been used together with linear regression coefficients fitted to measurement data. In fact, the 3G UMTS specification [1] includes multiple shifted, scaled, and quantized versions of CM depending on the system configuration. For 4G LTE, the CM fitting was updated [12] based on OFDM signal measurements, and different regression coefficients were recommended for two different bandwidth configurations. Since then, 5G NR has reused these definitions from [13], originally derived for 4G, and in some cases from 3G work, see e.g., [6]. However, applying the early-4G linear fit often leads to negative power de-rating estimates when evaluating modern low-PAPR waveforms, such as DFT-s-OFDM with $\pi/2$-BPSK and FDSS as well as other schemes, see e.g. [33], which raises questions about the practical meaning of these CM-based estimates.
@@
It is worth emphasizing that CM may provide a much better power de-rating estimation than PAPR, primarily because it is always used with fitted regression coefficients from measurements, whereas PAPR is typically evaluated in its raw form, without linear fitting. However, nothing prevents applying a linear regression to PAPR as well. In fact, the 3GPP contribution introducing CM [13] also considered a linear fitting of PAPR, $\approx 0.855({\rm PAPR}-3)$, showing that power de-rating correlates reasonably well with PAPR after regression, although CM provided a slightly better overall fit. More recent evaluations, such as [36], confirms this trend.
In any case, a single-regression of neither CM nor PAPR could fully predict realistic power de-rating in all scenarios. In particular, both metrics in their raw forms fail to capture the impact of bandwidth size and allocation position [8].
@@ IV-A PAPR Bounds
Moreover, for regular QAM, note that the maximum symbol power $A_{\mathcal{C}}^{2}$ occurs only for a subset of constellation points at the four edges, forming a QPSK sub-constellation with larger amplitude. We can therefore expect that symbol combination of this sub-QPSK constellation will yield some of the highest signal peaks, leading to:
${\rm PAPR}^{\rm U}(\mathcal{C}_{\rm QAM},N_{\rm e},L)\approx A_{\mathcal{C}_{\rm QAM}}^{2}\times{\rm PAPR}^{\rm U}(\mathcal{C}_{\rm QPSK},N_{\rm e},L).$ (16)
@@ IV-B Best Shifting Value
In Fig. 7, we compare the PAPR of QPSK with different values of $L$ as a function of the SE ratio $N_{\rm e}/N_{\rm sc}$, using the $10^{-3}$-CCDF PAPR with $N_{\rm sc}=96$ and Kaiser FDSS window with $\kappa=2$ (maximum power attenuation of $7.15$ dB). The PAPR with $L$ as in (18) and $L$ as in (19) are compared by deviation from the one with $L=0$. The figure also show a prediction of this PAPR deviation using the bound (13) with the corresponding $L$ values. As seen, $L$ as in (18), which corresponds to the symmetric extension of [31], when used with QPSK is always worse than $L=0$, which corresponds to the single-side extension in [35]. Similar results can be obtained for higher-order QAMs. The maximum PAPR gain over [31] by using a different $L$ is only a fraction of a dB, here up to 0.5 dB, and this gain decreases as the SE size increases.
@@ IV-C Optimum SE Size $N_{\rm e}^{({\rm min\,PAPR})}$ for Minimum PAPR
IV-C1 Observations on the Existence of an Optimal SE Size
Intuitive Explanation
Consider the pulse-shaping filters in (6), derived from the band-limited FDSS window $W[k]$ in (7). Their amplitude essentially follows a sinc shape, with more or less attenuated side lobes. The kernel filter shape in (7) is independent of $N_{\rm e}$, except for a scaling factor $\frac{1}{\sqrt{N_{\rm sc}-N_{\rm e}}}$, which causes the peak energy of each pulse to increase as $N_{\rm e}$ increases. In parallel, since there are $(N_{\rm sc}-N_{\rm e})$ pulses per OFDM symbol, the pulses are spaced $\frac{N_{\rm fft}}{N_{\rm sc}-N_{\rm e}}$ samples apart in (6). This spacing increases with $N_{\rm e}$, which in turn changes the degree of pulse overlap due the sinusoidal shape of the pulses. Both effects combined, depending of $N_{\rm e}$, their overall combining may decrease or increase as a function of $N_{\rm e}$, leading to a non-monotonic PAPR variation.
@@ IV-C1 Observations on the Existence of an Optimal SE Size
The existence of this global optimum SE size that minimizes the PAPR can be systematically verified by simulations across many scenarios. In Fig. 9, we plot several examples of PAPR versus the SE ratio $N_{\rm e}/N_{\rm sc}$. Each plot shows the maximum PAPR from $10^{6}$ Monte Carlo points, the $10^{-1}$, $10^{-2}$, and $10^{-3}$ CCDF, and the bounds from (13) and (15). All subplots is with QPSK, except Fig. 9(d) with $\pi/2$-BPSK and Fig. 9(e) with 16QAM. All have $N_{\rm sc}=96$ except Fig. 9(f) with $N_{\rm sc}=48$. The first row, i.e. Figs. 9(a)–(c), uses different FDSS windows, while the second row, i.e. Figs. 9(d)–(f), varies the modulation or bandwidth, all with a Kaiser FDSS window ($\kappa=2$). Shift values match single-side extension ($L=0$) in the literature, except Fig. 9(d) where $L$ is chosen according to (18).
@@
Fig. 9 clearly shows that the PAPR does not decrease monotonically with $N_{\rm e}$. In all cases, the PAPR decreases up to an optimum SE size $N_{\rm e}^{({\rm min\,PAPR})}$, beyond which it increases again. Selecting $N_{\rm e}>N_{\rm e}^{({\rm min\,PAPR})}$ is detrimental not only for spectral efficiency but also for PAPR reduction. Remark that even without FDSS [Fig. 9(a)], SE can provide some PAPR reduction and so $N_{\rm e}^{({\rm min\,PAPR})}>0$.
@@
From Figs. 9(a)–(c) one can note that $N_{\rm e}^{({\rm min\,PAPR})}$ changes with the FDSS window, while the PAPR curves are almost identical between QPSK [Fig. 9(b)] and 16-QAM [Fig. 9(e)]. Similarly, the PAPR behavior seems consistent for QPSK with different bandwidths [Figs. 9(b) and 9(f)].
IV-C2 Semi-Analytical Characterizations of $N_{\rm e}^{({\rm min\,PAPR})}$ from Bounds
Fig. 9 shows good agreement in the overall trend between the simulated PAPR and the PAPR bounds (13) and (15) as a function of $N_{\rm e}$. Accordingly, these bounds can be used to identify $N_{\rm e}^{({\rm min\,PAPR})}$. While an analytical determination of the bound minimizer appears intractable, its numerical evaluation is significantly less complex than direct PAPR simulations. This, in turn, facilitates the study of the optimum SE as a function of other system parameters.